Number of pizzas sold

Size Mon. Tues. Wed. Thu. Fri. Sat. Sun.
Small 28 35 28 36 37 34 45
Mediu
m

36 37 47 32 36 38 47
Large 41 42 40 35 51 56 61
Family 30 33 39 33 78 76 83

a) Write the information in the table above as a 4 × 7 matrix, P.

b) Write down a suitable matrix, A, that when P is multiplied by A the resulting matrix will give the
number of each size of pizza sold in the week.

c) Calculate the matrix PA.


d) Write down a suitable matrix, B, that when B is multiplied by P the resulting matrix will give the
number of pizzas sold on each day of the week.

e) Calculate the matrix BP.
1
Expert's answer
2020-05-25T19:12:53-0400

a) "P = \begin{pmatrix}
 28 &35& 28 &36 &37 &34 &45 \\
 36 &37 &47 &32 &36& 38& 47\\
41 &42 &40 &35 &51 &56 &61\\
30 &33& 39 &33& 78& 76& 83
\end{pmatrix}"


b) "A=\begin{pmatrix}
 1 \\
 1\\
1\\
1\\
1\\
1\\
1
\end{pmatrix}"

c) "PA=\begin{pmatrix}
 PA_{11} \\
 PA_{21}\\
PA_{31}\\
PA_{41}
\end{pmatrix}"


PA11 = P11 * A11 + P12 * A21 + P13 * A31 + P14 * A41 + P15 * A51 + P16 * A61 + P17 * A71 =

= 28 * 1 + 35 * 1 + 28 * 1 + 36 * 1 + 37 * 1 + 34 * 1 + 45 * 1 = 243

PA21 = P21 * A11 + P22 * A21 + P23 * A31 + P24 * A41 + P25 * A51 + P26 * A61 + P27 * A71 =

= 36 * 1 + 37 * 1 + 47 * 1 + 32 * 1 + 36 * 1 + 38 * 1 + 47 * 1 = 273

PA31 = P31 * A11 + P32 * A21 + P33 * A31 + P34 * A41 + P35 * A51 + P36 * A61 + P37 * A71 =

= 41 * 1 + 42 * 1 + 40 * 1 + 35 * 1 + 51 * 1 + 56 * 1 + 61 * 1 = 326

PA41 = P41 * A11 + P42 * A21 + P43 * A31 + P44 * A41 + P45 * A51 + P46 * A61 + P47 * A71 =

= 30 * 1 + 33 * 1 + 39 * 1 + 33 * 1 + 78 * 1 + 76 * 1 + 83 * 1 = 372


"PA=\begin{pmatrix}
 243 \\
 273\\
326\\
372
\end{pmatrix}"


d) "B=\begin{pmatrix}
 1 & 1 & 1 &1
\end{pmatrix}"

e) BP = (BP11 BP12 BP13 BP14 BP15 BP16 BP17)

BP11 = B11 * P11 + B12 * P21 + B13 * P31 + B14 * P41 = 1 * 28 + 1 * 36 + 1 * 41 + 1 * 30 = 135

BP12 = B11 * P12 + B12 * P22 + B13 * P32 + B14 * P42 = 1 * 35 + 1 * 37 + 1 * 42 + 1 * 33 = 147

BP13 = B11 * P13 + B12 * P23 + B13 * P33 + B14 * P43 = 1 * 28 + 1 * 47 + 1 * 40 + 1 * 39 = 154

BP14 = B11 * P14 + B12 * P24 + B13 * P34 + B14 * P44 = 1 * 36 + 1 * 32 + 1 * 35 + 1 * 33 = 136

BP15 = B11 * P15 + B12 * P25 + B13 * P35 + B14 * P45 = 1 * 37 + 1 * 36 + 1 * 51 + 1 * 78 = 202

BP16 = B11 * P16 + B12 * P26 + B13 * P36 + B14 * P46 = 1 * 34 + 1 * 38 + 1 * 56 + 1 * 76 = 204

BP17 = B11 * P17 + B12 * P27 + B13 * P37 + B14 * P47 = 1 * 45 + 1 * 47 + 1 * 61 + 1 * 83 = 236


BP = (135 147 154 136 202 204 236)